# DEAR: Depth-Enhanced Action Recognition

### v1

@@ 3 Experiments
Implementation Details. To assess the performance of our approach, we used an input video clip with 8 frames each of size $224\times 224$, and frames were chosen using static stride. The encoder that we are using is the ViT-B/16 architecture [7], for which we used AdamW optimizer. [22]
@@
Proposed RGB+Depth+ Depth(VideoMamba) ViT-B Clip-400M 70.8
Results Table 1 presents the overall results and comparison with other methods. Our framework outperforms recent works in top-1 accuracy. Our full model RGB+Depth+Depth(VideoMamba) achieves 70.8% accuracy; this is 4.2% and 0.6% greater than the individual data models VideoMamba-S and S4V that our approach is inspired by. The table also shows the fusion of depth and RGB, RGB+Depth, is 0.5% lower than when we fuse the mamba logits too.
We also examined the impact of depth maps on various classes and found significant improvements in specific activities. For example, in class 145, Stuffing something into something, the depth map helped it go up from 63% to 73%. Also, classes 12 and 98, Dropping something onto something and Pushing something so that it almost falls off but doesn’t, achieved a 5% improvement each by introducing depth maps. However, some classes, like 122, Rolling something on a flat surface, experienced a reduction of accuracy by 8%, likely due to the lack of depth needed by this class for classification.

### v2

@@ 3 Experiments
Implementation Details. To assess the performance of our approach, we used an input video clip with 8 frames each of size $224\times 224$, and frames were chosen using static stride. The encoder that we are using is the ViT-B/16 architecture [7], for which we used AdamW optimizer [22]. Also, we set the batch size during inference to 26 as the optimum value.
@@
Proposed RGB+Depth+ Depth(VideoMamba) ViT-B Clip-400M 71.0
Results Table 1 presents the overall results and comparison with other methods. Our framework outperforms recent works in top-1 accuracy. Our full model RGB+Depth+Depth(VideoMamba) achieves 71.0% accuracy; this is 4.4% and 0.8% greater than the individual data models VideoMamba-S and S4V that our approach is inspired by. The table also shows the fusion of depth and RGB, RGB+Depth, is 0.7% lower than when we fuse the mamba logits too.
We also examined the impact of depth maps on various classes and found significant improvements in specific activities. For example, in class 145, Stuffing something into something, the depth map helped it go up from 61.7% to 72.9%. Also, classes 65, 106, and 100, Pretending to be tearing something that is not tearable and Putting something into something, and Pushing something so that it slightly moves achieved 5.6%, and 7.5%, and 7.1% improvement, respectively, each by introducing depth maps. However, some classes, like 150, Tearing something just a little bit, experienced a reduction of accuracy by 3.2%, likely due to the lack of depth needed by this class for classification.